OpenStudy (anonymous):

please help!!!! x^2/x^2+3*x+14/x...............multiply and simplify to lowest terms

jimthompson5910 (jim_thompson5910):

notice how there is an x^2 term in the first numerator and an 'x' term in the second denominator

jimthompson5910 (jim_thompson5910):

so x^2 divided by x = x

jimthompson5910 (jim_thompson5910):

which means that \[\Large \frac{x^2}{x^2+3}*\frac{x+14}{x}\] becomes \[\Large \frac{x}{x^2+3}*\frac{x+14}{1}\]

jimthompson5910 (jim_thompson5910):

From here, you multiply the numerators together to get a final numerator then you multiply the denominators together to get a final denominator

OpenStudy (anonymous):

@jim_thompson5910, thanks for your help, so i came up with 14x^2/4x^2???

jimthompson5910 (jim_thompson5910):

hmm that's not correct

OpenStudy (anonymous):

or should it have been 14x^2/3x^3

jimthompson5910 (jim_thompson5910):

You should get this \[\Large \frac{x^2}{x^2+3}*\frac{x+14}{x}\] \[\Large \frac{x}{x^2+3}*\frac{x+14}{1}\] \[\Large \frac{x(x+14)}{(x^2+3)*1}\] \[\Large \frac{x^2+14x}{x^2+3}\]

OpenStudy (anonymous):

oh ok I see, what I done wrong

jimthompson5910 (jim_thompson5910):

glad you do

OpenStudy (anonymous):

thanks... is division pretty much the same way??

jimthompson5910 (jim_thompson5910):

with division, you flip the second fraction and you multiply

OpenStudy (anonymous):

ok can you just look at this one and tell me if Im way off or not??

jimthompson5910 (jim_thompson5910):

sure go ahead and post it

OpenStudy (anonymous):

5x^2-15x/2x-9 divided by x-3/2x-9...sorry could get the equations to work but I came up with 5x^2-15/x-3

jimthompson5910 (jim_thompson5910):

you should have 5x^2-15x/x-3 Then you factor the numerator: 5x^2 - 15x = 5x(x-3) The 'x-3' terms will cancel leaving you with just 5x as the final answer

OpenStudy (anonymous):

okay I get it..thank you again

jimthompson5910 (jim_thompson5910):

yw